How many metric tons (t) of oxygen are in a sample of pitchblende that contains 2.3 t of uranium?
(Hint: Sample of 84.2 g pitchblende contains 71.4 g of uranium and oxygen is the only other element present.)
1
Expert's answer
2018-03-14T08:47:47-0400
Uranium present in oxide with formula U2O8.

Ratio of oxygen in oxide=(8*16)/(238*2+8*16)
Ratio of oxygen in pitchblende =(8*16)/(238*2+8*16)*(71.4g/82.4g)
mass of oxygen =(8*16)/(238*2+8*16)*(71.4g/82.4g)*2.3t=0.45t